Q.(((The sum of two times one number and three times another is 29. The difference between twice the first and the second is 1. Solve algebraically to find the numbers))).
Answer by stanbon(75887)

The sum of two times one number and three times another is 29. The difference between twice the first and the second is 1. Solve algebraically to find the numbers
-------------------------------
Equations:
2x + 3y = 29
2x - y = 1
-------------------
Subtract 2nd from 1st to solve for "y":
4y = 28
y = 7
-------------------
Since 2x - y = 1, 2x -7 = 1; 2x = 8; x = 4
------------------------------------------------
Ans: 1st = x = 4 ; 2nd = y = 7
==================================================
Cheers,
Stan H.